OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a^2 - 3b^3 + c^2 + 2b^3 - c^2 if a = 3, b = 2, and c = -3.

OpenStudy (anonymous):

-3 x -3 = 9

OpenStudy (anonymous):

Whoops hold on forgot to put these up

OpenStudy (anonymous):

-3^2 - 3(2)^3 + -3^2 + 2(2)^3 - -3^2

OpenStudy (anonymous):

All you have to do is substitute a 3 where there is an a, a 2 where there is a b, and a -3 where there is a c. It looks like you're doing this right \[-a ^{2}-3b ^{3}+c ^{2}+2b ^{3}-c ^{2}\]

OpenStudy (anonymous):

Yea but I'm coming up with the wrong answer :\ so I'm not really sure.

OpenStudy (anonymous):

Hold on let me try to solve it

OpenStudy (anonymous):

ok thanks lol

OpenStudy (anonymous):

\[-(3)^{2}-3(2)^{3}+(-3)^{2}+2(2)^{3}-(-3)\]

OpenStudy (anonymous):

31

OpenStudy (anonymous):

-31*

OpenStudy (anonymous):

Oops I made a typo The last one should be ^2 It would look like that right? Then just simplify -3 * -3 = 9 - 3 * 2 = -5, -5 * -5 * -5 = -125 -3 * -3 = 9 2 * 2 = 4, 4 * 4 * 4 = 64 3 * 3 = 9

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (anonymous):

I got -31

OpenStudy (anonymous):

That's what I got too :)

OpenStudy (anonymous):

:) haha thanks for taking your time with me, wasn't really sure.

OpenStudy (anonymous):

Yw :p